OpenStudy (anonymous):

Determine whether the sequence below is a geometric sequence and, if so, find a formula that describes the sequence. 1, 3, 9, 27, 81, …

OpenStudy (anonymous):

here 3/1 =9/3 =27/9 and so on so it is a geometric sequence its nth term is given by 3^(n-1) n>=1 and n belongs to N